the perimeter of rectangle is 66 inches. Find the dimensions its length is 1 more than 3 times its width.

Answer by bluemockingjay7(41)

perimeter = 66 inches
length = 3w + 1
width = w
3w + 1 + w = 66
2(3w + 1 + w)= 66 (since 2(L + W)= P)
6w + 2 + 2w = 66
8w + 2 = 66
8w = 66 - 2
8w = 64
8w/8 = 64/8
w = 8
width = 8 inches
length = 3(8) + 1 = 24 + 1 = 25 inches